OpenStudy (wade123):

Find an equation for the nth term of the sequence. -1, -4, -16, -64, ...

OpenStudy (anonymous):

n x 4

OpenStudy (wade123):

my answer

OpenStudy (ikram002p):

-\( 4^n\)

OpenStudy (wade123):

options:

OpenStudy (ikram002p):

ur answer is wrong :O

OpenStudy (wade123):

nooo :(

OpenStudy (anonymous):

i would say its c..

OpenStudy (ikram002p):

does n start from 0 ? if yes then its -1 . 4^n if n starts from 1 then its -1. 4^(n-1)

OpenStudy (ikram002p):

@ninjasandtigers C is wrong answer

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

so @wade123 does n start from 0 or 1 ?

OpenStudy (wade123):

sorry my internet is really slow right now @ikram002p

OpenStudy (wade123):

1

OpenStudy (ikram002p):

from 0 ?

OpenStudy (wade123):

-1

OpenStudy (ikram002p):

?

OpenStudy (ikram002p):

xD

OpenStudy (ikram002p):

i told yo , if n start from 0 then its -1 . 4^n if n start from 1 then its -1.4^(n-1) @ganeshie8 is it comon that thes things start from 1 ?

OpenStudy (wade123):

im so confused.. @ikram002p

OpenStudy (ikram002p):

im confused like u :-\ let \(a_n=-1*4^n\) then \( a_0=-1\) \(a_1=-4\) \(a_2=-16\) right ? let \(\large a_n=-1 *4^{n-1}\) then \( a_1=-1\) \(a_2=-4\) \(a_3=-16\)

OpenStudy (wade123):

so it would be the first one? @ikram002p

OpenStudy (ikram002p):

in ur text book does n in sequenc always start from 0 or 1 ?

OpenStudy (wade123):

@ikram002p you wrote the same exact thing for each!

OpenStudy (wade123):

@ganeshie8 please help

ganeshie8 (ganeshie8):

if n starts from 1 then its -1. 4^(n-1)

ganeshie8 (ganeshie8):

\(\large a_n = -1 \bullet 4^{n-1}\)

ganeshie8 (ganeshie8):

in textbook, usually n starts from 1

OpenStudy (ikram002p):

xD

OpenStudy (wade123):

im a little confused but i think i got it

OpenStudy (wade123):

so thats it right @ganeshie8 ? I cant tell if youre still typing